find the standard form of the equation of the circle with the endpoints of the diameter at the points of (5,2) and (-9,6)
Answer by Alan3354(69443)

find the standard form of the equation of the circle with the endpoints of the diameter at the points of (5,2) and (-9,6)
-----------
The diameter is the distance between the 2 points.
d^2 = diffy^2 + diffx^2 = 4^2 + 14^2
d^2 = 212
-----------
The center is the midpoint. Find the average of x and y separately.
x: (5-9)/2 = -2
y: (2+6)/2 = 4
Center at (2,-4)
(x-2)^2 + (y-4)^2 = 212 is the circle